# A Novel Hybrid Deep Learning Technique for Speech Emotion Detection using Feature Engineering

### v1

@@ Abstract
HHAI
@@ 1 Introduction
With technological advancements, research in human-computer interaction (HCI) and artificial emotional intelligence (AEI) is evolving rapidly [1, 4]. HCI [5] focuses on how people communicate with computers and how effectively systems respond. Researchers aim to make interactions between humans and machines as natural as possible. Speech, being our primary form of communication, plays a key role in HCI, with microphones as auditory sensors [6]. Through speech, we express emotions, making speech emotion recognition (SER) essential for enhancing HCI [7].
Speech emotion detection (SED) or SER is a sub-domain of natural language processing (NLP) [8, 2] and affective computing [9]. There has been a significant amount of published work in recent years [10, 11]. A conventional SER system uses various methods to extract and analyze speech signals to identify emotions [11]. SER has many practical applications, especially in enhancing human-computer interactions by adding emotional awareness [12].
One practical example of SER includes assessing call center agents’ performance by detecting customer emotions such as anger or happiness. This feedback helps companies improve service quality, offer targeted training, and increase customer satisfaction and efficiency [13]. Beyond this, many applications—including healthcare, smart home devices, audio surveillance, criminal investigations, recommendation systems, dialogue systems, and intelligent robots—benefit from detecting users’ emotions through speech.
Despite advances, challenges in SER remain due to limited technology, the complex nature of emotions, and variations in language, accents, gender, and age, which all impact how emotions are expressed in speech [14, 15]. To address these limitations, researchers have turned to cross-dataset integration techniques [16, 17]. By combining diverse datasets, models can learn more generalized patterns and improve recognition accuracy across multiple contexts [18].
@@ 2 Related Work
SER identifies speakers’ emotional states from speech and plays a vital role in enhancing HCI across domains like call centers, mental health diagnostics, and driver monitoring systems. By interpreting emotional cues, SER improves user interaction, enables emotionally adaptive interfaces, and supports stress or anxiety detection [19, 20].
@@
The traditional approach to SER has primarily emphasized extracting prosodic features—such as pitch, intensity, and energy—and spectral features like Mel-frequency cepstral coefficients (MFCC) and Linear Predictive Coefficients (LPC), which are essential for representing emotional cues in speech [10]. These features have typically been evaluated using machine learning classifiers such as Gaussian Mixture Models and Support Vector Machines (SVM) [21, 20]. While earlier studies focused on a limited set of handcrafted features for model development [20], there remains limited research exploring a broader range of acoustic feature types. In contrast, our work leverages a more comprehensive and diverse set of features—including MFCC, Chroma, Spectral Contrast, RMSE, ZCR, and Log-Mel Spectrograms—resulting in a high-dimensional feature space. This extensive variation not only captures more nuanced emotional characteristics but may also contribute to improved classification performance compared to prior works that relied on fewer features.
@@
Deep learning has revolutionized various fields, including SER, by providing advanced methods for processing and analyzing complex data. Techniques such as Convolutional Neural Networks (CNNs), Recurrent Neural Networks (RNNs), Long Short-Term Memory networks (LSTMs), and attention mechanisms have significantly enhanced the accuracy and efficiency of SER systems. In SER, CNNs have been used to extract high-level features from spectrograms, which are then used for emotion classification [20]. RNNs are designed to process sequential data, making them suitable for time-series analysis, such as speech signals [11]. LSTMs, a type of RNN, address the vanishing gradient problem and are capable of learning long-term dependencies, which is beneficial for SER [18]. Combining CNNs with LSTMs allows for the extraction of both spatial and temporal features, enhancing the performance of SER systems [22, 11]. Deep learning has significantly improved the accuracy and robustness of SER systems by enabling the automatic extraction and processing of complex features from raw data [20].
@@
The study of emotion recognition using speech datasets is a key area in AI and machine learning. Datasets like RAVDESS, TESS, SAVEE, EMO-DB, and CREMA-D provide diverse audio samples for training and evaluation. RAVDESS offers high-quality recordings and wide emotion coverage, achieving 97.16% accuracy with ConvLSTM [23], though its limited size may cause overfitting [24]. TESS has 2,800 samples and shows perfect accuracy in some studies [23], but lacks speaker diversity [3]. SAVEE, focused on male British speech, reached 97.45% accuracy [3, 23], yet its small size is a limitation [24, 23]. EMO-DB supports cross-linguistic studies [23] with 535 German samples but faces overfitting risks [24, 23]. CREMA-D includes varied demographics [3], though its complexity impacts consistent training despite achieving 83.28% in some works [23]. While valuable, these datasets highlight the need for more diverse and extensive data to improve SER model generalization.
@@
Despite notable progress in deep learning-based SER, key challenges persist in achieving high accuracy and robust generalization across diverse datasets. Variations in emotional expression across individuals, languages, and cultural backgrounds make it difficult for models to generalize effectively [23]. The inherent subjectivity of emotions also introduces ambiguity in labeling, leading to inconsistencies in classification outcomes [21]. Moreover, many existing SER datasets are scripted and limited in speaker and emotional diversity, which reduces their applicability to real-world scenarios and increases the risk of overfitting [3]. Issues such as class imbalance and small dataset sizes further hinder deep learning model performance and stability [3]. While deep models like CNNs and LSTMs can automatically learn features, selecting relevant and informative features remains a challenge [21]. Additionally, the computational demands of these architectures can limit their use in real-time systems. Although techniques like data augmentation and cross-dataset integration offer potential improvements by enhancing training diversity [23, 3], relatively few studies explore the impact of high-dimensional and varied acoustic features in conjunction with structured prediction models. These limitations motivate our work, which addresses the need for better feature diversity, improved generalization, and context-aware modeling in SER.
@@ 3.1 Dataset
In this study, we used five publicly available datasets namely Ryerson Audio-Visual Database of Emotional Speech and Song (RAVDESS), Toronto Emotional Speech Set (TESS), Surrey Audio-Visual Expressed Emotion (SAVEE), EmoDB and CREMA-D, and for speech emotion detection. Table 1 shows the emotion count of each of the five datasets and combined datasets class-wise [3]. All datasets used in this study are publicly available and widely used for academic research. We did not collect any new human data. Dataset diversity, spanning gender, age, and language, was considered to enhance fairness and generalizability.
@@
  • Ryerson Audio-Visual Database of Emotional Speech and Song (RAVDESS): This dataset[25] consists of 1440 audio wav files recorded by 12 males and 12 females. It includes emotions such as happy, calm, sad, fearful, angry, surprise, and disgust. The emotions are recorded at normal and strong intensity levels. It is a balanced dataset though the "neutral" class has fewer records compared to other classes. We exclude the calm emotion class from RAvDESS to matched with the combination dataset.
  • Toronto Emotional Speech Set (TESS): A female-only dataset [26] consisting of 2800 audio files recorded by two actors aged 26 and 64 years. The recordings portray seven emotions: happiness, anger, fear, disgust, pleasant surprise, sadness, and neutral.
  • Surrey Audio-Visual Expressed Emotion (SAVEE): This dataset [27] consists of 120 audio clips recorded by four male speakers in the age group of 27 to 31 years. The emotions included are neutral, happy, sad, angry, disgust, fear and surprise. However, this dataset has a class imbalance issue, with the "neutral" class is almost double compared to all the other classes.
  • EmoDB: It [28] comprises 535 audio recordings in the German language categorized into seven emotional kinds: "anger," "fear," "sadness," "happiness," "disgust," "boredom," and "neutral". The utterances are sampled at a rate of 16 kHz with a resolution of 16 bits. However, this dataset has a class imbalance issue, with the "anger," class utterance number is large compared to other classes. We exclude the boredom emotion class from EmoDB to matched with the combination dataset.
  • CREMAD: The CREMA-D [29] dataset contains 7,442 audio files recorded by 91 actors (48 males and 43 females) between the age group of 20 to 74 years. It consists of six emotions, namely happy, angry, neutral, sad, fear and disgust at four different emotions levels (low, medium, high, and unspecified)
@@ 3.1.2 Data Augmentation
To create variations in audio data, we applied the data augmentation process [30]. Data augmentation is a widely used technique in SER model to enhance the dataset, improve robustness and generalization. It’s also increased data size for proper training of the models. In this study, we noticed the data set imbalance over the emotion classes, so we injected Gaussian noise into three faces. Our noise injection technique applied with the noise rate (0.035, 0.025, 0.015). We also applied two more augmentation techniques to our dataset, pitch shift and time stretch, by using librosa. The pitch shift technique shifts the pitch of an audio signal without changing its duration in three steps with the rate parameter of (0.70, 0.80, 0.70). And we applied the time stretch technique to change the speed of an audio signal without altering its pitch with the speed factor rate (0.8, 0.9, 0.7). These techniques enhance the model’s ability to handle different acoustic environments, improving the overall performance of the SER model. The impact of these augmentation techniques is visually shown in Figure 4 . Spectrogram and waveform visualizations illustrating the effects of different data augmentation techniques applied to a speech sample, Time-stretching increases the playback speed while preserving pitch, Pitch shifting alters the frequency content to simulate different speaker tones, and Gaussian noise injection simulates environmental noise to enhance model robustness.
@@ 3.2 Feature Extraction and Selection
  • MFCCs (Mel-Frequency Cepstral Coefficients) : The sounds produced by humans are shaped by the unique configuration of the vocal tract, including features such as the tongue and teeth. These structures influence each individual’s distinct voice characteristics. Accurately capturing vocal tract shape is important, as it is reflected in the short-time power spectrum, typically represented by Mel-Frequency Cepstral Coefficients (MFCCs) [22]. MFCCs are widely used in SER research [31, 3, 32]. To extract MFCCs, the speech signal is segmented into overlapping frames of 20–30 ms with a 10 ms shift to preserve temporal dynamics. Each frame is windowed and transformed using the Discrete Fourier Transform (DFT) to obtain magnitude spectrum [19, 33]. A set of 26 Mel-scaled filters is applied to mimic the human auditory system, producing 26 energy values per frame. These are converted into log filter bank energies. The relationship between Mel and physical frequency is expressed by Equation 1 .
@@
  • Chroma Features: The Chroma feature captures the tonal content of audio by mapping frequency components to the 12 pitch classes of the chromatic scale [34]. It reflects energy distribution across pitch classes as a 12-dimensional vector. Chroma features are extracted using Short-Time Fourier Transform (STFT) on audio waveforms [35]. This study used three Chroma types: Chroma-STFT, Chroma-CQT (Constant-Q Transform), and Chroma-CENS (Chroma Energy Normalized Statistics). Each captures pitch information differently. From each audio file, 12 features were extracted per type, totaling 36 Chroma-related features. These represent tonal and harmonic speech content, aiding emotion recognition.
  • Log Mel Spectrogram (LMS): The Log Mel spectrogram improves audio spectrum representation by considering temporal changes and the human ear’s frequency response. It includes signal segmentation, windowing, the Fourier transform, Mel filtering, and log compression to generate a logarithmic spectral view. The frequency spectrum is mapped onto Mel scale frequencies, forming a Mel spectrogram per window. Magnitude components of these frequencies were extracted using the Librosa library. The spectrogram visualizes signal intensity in the time-frequency domain using the Fast Fourier Transform (FFT) [20, 36]. It is essential for speech classification tasks and is effective with BiLSTM or DeepCRF. In this study, 64 LMS features were extracted from each audio file.
  • Spectral Contrast: Spectral entropy serves as an indicator of voicing and signal quality [37]. It is often utilized to assess the ’speechiness’ of an audio signal and is widely applied in distinguishing speech from background noise.
  • Root Mean Square Energy (RMSE): The Root Mean Square Energy gives the average signal amplitude over a frame, regardless of whether the values are positive or negative. It is useful for analyzing signal intensity in speech [21]. In this study, RMSE values were extracted using the Librosa library. For a signal $x=x_{1},x_{2},...,x_{n}$, the RMSE is calculated using Equation 2 . In this study, we extracted 3 features for RMSE.
@@
  • Zero Crossing Rate (ZCR): Zero Crossing Rate is a commonly used feature in SER. It counts the number of times a signal crosses the zero value in a given time frame. This helps to differentiate between voiced and unvoiced parts of speech [3]. In this work, we used the Librosa library to extract ZCR values from the audio datasets. The ZCR is defined mathematically in Equation 3

### v2

@@ Abstract
HHAI-WS 2025: Workshops at the Fourth International Conference on Hybrid Human-Artificial Intelligence (HHAI), June 9–13, 2025, Pisa, Italy
[…]
Copyright for this paper by its authors. Use permitted under Creative Commons License Attribution 4.0 International (CC BY 4.0).
@@ 1 Introduction
With technological advancements, research in human-computer interaction (HCI) and artificial emotional intelligence (AEI) is evolving rapidly [krakovsky2018artificial, schuller2018age]. HCI [hudlicka2003feel] focuses on how people communicate with computers and how effectively systems respond. Researchers aim to make interactions between humans and machines as natural as possible. Speech, being our primary form of communication, plays a key role in HCI, with microphones as auditory sensors [mustaqeem2019cnn]. Through speech, we express emotions, making speech emotion recognition (SER) essential for enhancing HCI [chatterjee2018speech].
Speech emotion detection (SED) or SER is a sub-domain of natural language processing (NLP) [brown2020language, kusal2023systematic] and affective computing [picard2000affective]. There has been a significant amount of published work in recent years [madanian2023speech, akccay2020speech]. A conventional SER system uses various methods to extract and analyze speech signals to identify emotions [akccay2020speech]. SER has many practical applications, especially in enhancing human-computer interactions by adding emotional awareness [mustafa2018speech].
One practical example of SER includes assessing call center agents’ performance by detecting customer emotions such as anger or happiness. This feedback helps companies improve service quality, offer targeted training, and increase customer satisfaction and efficiency [gadhe2015emotion]. Beyond this, many applications—including healthcare, smart home devices, audio surveillance, criminal investigations, recommendation systems, dialogue systems, and intelligent robots—benefit from detecting users’ emotions through speech.
Despite advances, challenges in SER remain due to limited technology, the complex nature of emotions, and variations in language, accents, gender, and age, which all impact how emotions are expressed in speech [radhika2025reliable, kowalczyk2012detecting]. To address these limitations, researchers have turned to cross-dataset integration techniques [ottoni2023deep, milner2019cross]. By combining diverse datasets, models can learn more generalized patterns and improve recognition accuracy across multiple contexts [jothimani2022mff].
@@ 2 Related Work
SER identifies speakers’ emotional states from speech and plays a vital role in enhancing HCI across domains like call centers, mental health diagnostics, and driver monitoring systems. By interpreting emotional cues, SER improves user interaction, enables emotionally adaptive interfaces, and supports stress or anxiety detection [nantasri2020light, mukhamediya2023effect].
@@
The traditional approach to SER has primarily emphasized extracting prosodic features—such as pitch, intensity, and energy—and spectral features like Mel-frequency cepstral coefficients (MFCC) and Linear Predictive Coefficients (LPC), which are essential for representing emotional cues in speech [madanian2023speech]. These features have typically been evaluated using machine learning classifiers such as Gaussian Mixture Models and Support Vector Machines (SVM) [er2020novel, mukhamediya2023effect]. While earlier studies focused on a limited set of handcrafted features for model development [mukhamediya2023effect], there remains limited research exploring a broader range of acoustic feature types. In contrast, our work leverages a more comprehensive and diverse set of features—including MFCC, Chroma, Spectral Contrast, RMSE, ZCR, and Log-Mel Spectrograms—resulting in a high-dimensional feature space. This extensive variation not only captures more nuanced emotional characteristics but may also contribute to improved classification performance compared to prior works that relied on fewer features.
@@
Deep learning has revolutionized various fields, including SER, by providing advanced methods for processing and analyzing complex data. Techniques such as Convolutional Neural Networks (CNNs), Recurrent Neural Networks (RNNs), Long Short-Term Memory networks (LSTMs), and attention mechanisms have significantly enhanced the accuracy and efficiency of SER systems. In SER, CNNs have been used to extract high-level features from spectrograms, which are then used for emotion classification [mukhamediya2023effect]. RNNs are designed to process sequential data, making them suitable for time-series analysis, such as speech signals [akccay2020speech]. LSTMs, a type of RNN, address the vanishing gradient problem and are capable of learning long-term dependencies, which is beneficial for SER [jothimani2022mff]. Combining CNNs with LSTMs allows for the extraction of both spatial and temporal features, enhancing the performance of SER systems [gupta2022detecting, akccay2020speech]. Deep learning has significantly improved the accuracy and robustness of SER systems by enabling the automatic extraction and processing of complex features from raw data [mukhamediya2023effect].
@@
The study of emotion recognition using speech datasets is a key area in AI and machine learning. Datasets like RAVDESS, TESS, SAVEE, EMO-DB, and CREMA-D provide diverse audio samples for training and evaluation. RAVDESS offers high-quality recordings and wide emotion coverage, achieving 97.16% accuracy with ConvLSTM [ben2024enhancing], though its limited size may cause overfitting [issa2020speech]. TESS has 2,800 samples and shows perfect accuracy in some studies [ben2024enhancing], but lacks speaker diversity [ahmed2023ensemble]. SAVEE, focused on male British speech, reached 97.45% accuracy [ahmed2023ensemble, ben2024enhancing], yet its small size is a limitation [issa2020speech, ben2024enhancing]. EMO-DB supports cross-linguistic studies [ben2024enhancing] with 535 German samples but faces overfitting risks [issa2020speech, ben2024enhancing]. CREMA-D includes varied demographics [ahmed2023ensemble], though its complexity impacts consistent training despite achieving 83.28% in some works [ben2024enhancing]. While valuable, these datasets highlight the need for more diverse and extensive data to improve SER model generalization.
@@
Despite notable progress in deep learning-based SER, key challenges persist in achieving high accuracy and robust generalization across diverse datasets. Variations in emotional expression across individuals, languages, and cultural backgrounds make it difficult for models to generalize effectively [ben2024enhancing]. The inherent subjectivity of emotions also introduces ambiguity in labeling, leading to inconsistencies in classification outcomes [er2020novel]. Moreover, many existing SER datasets are scripted and limited in speaker and emotional diversity, which reduces their applicability to real-world scenarios and increases the risk of overfitting [ahmed2023ensemble]. Issues such as class imbalance and small dataset sizes further hinder deep learning model performance and stability [ahmed2023ensemble]. While deep models like CNNs and LSTMs can automatically learn features, selecting relevant and informative features remains a challenge [er2020novel]. Additionally, the computational demands of these architectures can limit their use in real-time systems. Although techniques like data augmentation and cross-dataset integration offer potential improvements by enhancing training diversity [ben2024enhancing, ahmed2023ensemble], relatively few studies explore the impact of high-dimensional and varied acoustic features in conjunction with structured prediction models. These limitations motivate our work, which addresses the need for better feature diversity, improved generalization, and context-aware modeling in SER.
@@ 3.1 Dataset
In this study, we used five publicly available datasets namely Ryerson Audio-Visual Database of Emotional Speech and Song (RAVDESS), Toronto Emotional Speech Set (TESS), Surrey Audio-Visual Expressed Emotion (SAVEE), EmoDB and CREMA-D, and for speech emotion detection. Table 1 shows the emotion count of each of the five datasets and combined datasets class-wise [ahmed2023ensemble]. All datasets used in this study are publicly available and widely used for academic research. We did not collect any new human data. Dataset diversity, spanning gender, age, and language, was considered to enhance fairness and generalizability.
@@
  • Ryerson Audio-Visual Database of Emotional Speech and Song (RAVDESS): This dataset[livingstone2018ryerson] consists of 1440 audio wav files recorded by 12 males and 12 females. It includes emotions such as happy, calm, sad, fearful, angry, surprise, and disgust. The emotions are recorded at normal and strong intensity levels. It is a balanced dataset though the "neutral" class has fewer records compared to other classes. We exclude the calm emotion class from RAvDESS to matched with the combination dataset.
  • Toronto Emotional Speech Set (TESS): A female-only dataset [pichora2020toronto] consisting of 2800 audio files recorded by two actors aged 26 and 64 years. The recordings portray seven emotions: happiness, anger, fear, disgust, pleasant surprise, sadness, and neutral.
  • Surrey Audio-Visual Expressed Emotion (SAVEE): This dataset [jackson2014surrey] consists of 120 audio clips recorded by four male speakers in the age group of 27 to 31 years. The emotions included are neutral, happy, sad, angry, disgust, fear and surprise. However, this dataset has a class imbalance issue, with the "neutral" class is almost double compared to all the other classes.
  • EmoDB: It [burkhardt2005database] comprises 535 audio recordings in the German language categorized into seven emotional kinds: "anger," "fear," "sadness," "happiness," "disgust," "boredom," and "neutral". The utterances are sampled at a rate of 16 kHz with a resolution of 16 bits. However, this dataset has a class imbalance issue, with the "anger," class utterance number is large compared to other classes. We exclude the boredom emotion class from EmoDB to matched with the combination dataset.
  • CREMAD: The CREMA-D [cao2014crema] dataset contains 7,442 audio files recorded by 91 actors (48 males and 43 females) between the age group of 20 to 74 years. It consists of six emotions, namely happy, angry, neutral, sad, fear and disgust at four different emotions levels (low, medium, high, and unspecified)
@@ 3.1.2 Data Augmentation
To create variations in audio data, we applied the data augmentation process [ferreira2022survey]. Data augmentation is a widely used technique in SER model to enhance the dataset, improve robustness and generalization. It’s also increased data size for proper training of the models. In this study, we noticed the data set imbalance over the emotion classes, so we injected Gaussian noise into three faces. Our noise injection technique applied with the noise rate (0.035, 0.025, 0.015). We also applied two more augmentation techniques to our dataset, pitch shift and time stretch, by using librosa. The pitch shift technique shifts the pitch of an audio signal without changing its duration in three steps with the rate parameter of (0.70, 0.80, 0.70). And we applied the time stretch technique to change the speed of an audio signal without altering its pitch with the speed factor rate (0.8, 0.9, 0.7). These techniques enhance the model’s ability to handle different acoustic environments, improving the overall performance of the SER model. The impact of these augmentation techniques is visually shown in Figure 4 . Spectrogram and waveform visualizations illustrating the effects of different data augmentation techniques applied to a speech sample, Time-stretching increases the playback speed while preserving pitch, Pitch shifting alters the frequency content to simulate different speaker tones, and Gaussian noise injection simulates environmental noise to enhance model robustness.
@@ 3.2 Feature Extraction and Selection
  • MFCCs (Mel-Frequency Cepstral Coefficients) : The sounds produced by humans are shaped by the unique configuration of the vocal tract, including features such as the tongue and teeth. These structures influence each individual’s distinct voice characteristics. Accurately capturing vocal tract shape is important, as it is reflected in the short-time power spectrum, typically represented by Mel-Frequency Cepstral Coefficients (MFCCs) [gupta2022detecting]. MFCCs are widely used in SER research [de2023ongoing, ahmed2023ensemble, singh2023speech]. To extract MFCCs, the speech signal is segmented into overlapping frames of 20–30 ms with a 10 ms shift to preserve temporal dynamics. Each frame is windowed and transformed using the Discrete Fourier Transform (DFT) to obtain magnitude spectrum [nantasri2020light, hajarolasvadi20193d]. A set of 26 Mel-scaled filters is applied to mimic the human auditory system, producing 26 energy values per frame. These are converted into log filter bank energies. The relationship between Mel and physical frequency is expressed by Equation 1 .
@@
  • Chroma Features: The Chroma feature captures the tonal content of audio by mapping frequency components to the 12 pitch classes of the chromatic scale [birajdar2020speech]. It reflects energy distribution across pitch classes as a 12-dimensional vector. Chroma features are extracted using Short-Time Fourier Transform (STFT) on audio waveforms [kumar2021cnn]. This study used three Chroma types: Chroma-STFT, Chroma-CQT (Constant-Q Transform), and Chroma-CENS (Chroma Energy Normalized Statistics). Each captures pitch information differently. From each audio file, 12 features were extracted per type, totaling 36 Chroma-related features. These represent tonal and harmonic speech content, aiding emotion recognition.
  • Log Mel Spectrogram (LMS): The Log Mel spectrogram improves audio spectrum representation by considering temporal changes and the human ear’s frequency response. It includes signal segmentation, windowing, the Fourier transform, Mel filtering, and log compression to generate a logarithmic spectral view. The frequency spectrum is mapped onto Mel scale frequencies, forming a Mel spectrogram per window. Magnitude components of these frequencies were extracted using the Librosa library. The spectrogram visualizes signal intensity in the time-frequency domain using the Fast Fourier Transform (FFT) [mukhamediya2023effect, meng2019speech]. It is essential for speech classification tasks and is effective with BiLSTM or DeepCRF. In this study, 64 LMS features were extracted from each audio file.
  • Spectral Contrast: Spectral entropy serves as an indicator of voicing and signal quality [kumar2021analysis]. It is often utilized to assess the ’speechiness’ of an audio signal and is widely applied in distinguishing speech from background noise.
  • Root Mean Square Energy (RMSE): The Root Mean Square Energy gives the average signal amplitude over a frame, regardless of whether the values are positive or negative. It is useful for analyzing signal intensity in speech [er2020novel]. In this study, RMSE values were extracted using the Librosa library. For a signal $x=x_{1},x_{2},...,x_{n}$, the RMSE is calculated using Equation 2 . In this study, we extracted 3 features for RMSE.
@@
  • Zero Crossing Rate (ZCR): Zero Crossing Rate is a commonly used feature in SER. It counts the number of times a signal crosses the zero value in a given time frame. This helps to differentiate between voiced and unvoiced parts of speech [ahmed2023ensemble]. In this work, we used the Librosa library to extract ZCR values from the audio datasets. The ZCR is defined mathematically in Equation 3